OpenStudy (anonymous):

what does x equal if you use the quadratic formula? 2x^2+x-9=0

OpenStudy (callisto):

the quadratic equation having general form is ax^2 +bx +c =0 and the quadratic formula is\[x = (-b \pm \sqrt{b^2 - 4ac})/(2a)\] so put the value a=2, b=1 and c=-9 into the formula and see what you get

OpenStudy (callisto):

\[x = (-1 \pm \sqrt{1^2 - 4(2)(-9)})/[2(2)]\]